### v1

@@ Abstract
HIP 67522 is a young, 17 million year old star-planet system in the Upper Centaurus Lupus part of the Sco-Cen OB association [8, 5] about 125125125\,125pc away [9]. The star is a $1.2M_{\odot}$ dwarf [5] that hosts one gas giant planet in a 6.956.956.95\,6.95d orbit [5], and another in a 14.3314.3314.33\,14.33d orbit [6].
@@
HIP 67522 was previously observed by the Transiting Exoplanet Survey Satellite (TESS [15]) in Sectors 11 (May 2019), 38 (May 2021), and 64 (April 2023) at a 2-min cadence for a total of 65.665.665.6\,65.6d (Extended Data Fig. 5). Additionally, we gathered dedicated observations of the HIP 67522 system with the CHaracterising ExoPlanets Telescope (CHEOPS [16]) for a total of 7.67.67.6\,7.6d between March 9 and June 22, 2024 (Extended Data Table 1). We achieved excellent phase coverage: each hour of the 167167167\,167h orbit of HIP 67522  b was fully covered at least seven times (Fig. 3 and Extended Data Table 2). We detected 12 flares in the TESS light curves (Extended Data Fig. 6), and 4 in the CHEOPS light curves (Extended Data Fig. 7, see also Methods). Of the CHEOPS flares, 3 were above the detection threshold for TESS. We use these for analysis, giving a total of 15 flares (Extended Data Table 3), to ensure that the different detection thresholds for flares in CHEOPS and TESS do not bias our conclusion.
Figure 2: Observed flare rates on HIP 67522 relative to the orbit of the innermost planet HIP 67522  b. HIP 67522 showed a significant increase in flare rate in the expected range, indicating that the majority of detected flares were triggered by the planet. The orange dashed region illustrates the expected rate of planet-induced flares from a footpoint of interaction at the sub-planetary longitude. Planet-induced flares can only be seen when the planet faces the observer (±$\pm 90^{\circ}$), and most flares can be expected near the transit of HIP 67522  b ($0^{\circ}$).
@@
The best-fit center of the clustering region is at phase $0.08\pm 0.04$, consistent with being centered on transit within $2\sigma$. The slight offset from transit may therefore be coincidental. Alternatively, it may be due to a distinctive feature of the HIP 67522 system. With a rotation period of approximately 1.421.421.42\,1.42d, HIP 67522  b orbits near a 1:5 commensurability between the rotation period of the star and the orbital period of HIP 67522  b. If this resonance was exact, the planet could align with the same stellar longitude at fixed orbital phases. In this scenario, a misaligned stellar magnetic field (Fig. 1) would cause the planet to encounter periodic maxima and minima in magnetic field strength multiple times per orbit. This varying field induces harmonic modulation in the star-planet interaction strength, which depends on the position of the footpoint, and the stellar magnetic field strength encountered by the planet [1, 19]. This harmonic modulation introduces two parameters into the model: the dipole field obliquity $\theta$ and the initial dipole azimuthal orientation $\phi_{0}$ relative to the observer. These parameters reproduce the peak power of interaction in the observed phase range for about $17\%$ of all possible configurations, spanning all obliquities above $20^{\circ}$ (Fig. 4). If the system is not in resonance, this modulation would average out over time, leaving the clustering determined only by self-shadowing and visibility of the footpoint, similar to Fig. 2.
@@
Regardless of the underlying mechanism, the observed planet-induced emission has striking consequences. The orbital phases of the flares induced by HIP 67522  b imply that the interaction with the star backfires on the planet through self-inflicted space weather. A planet’s space weather denotes the bombardment with high energy radiation and particles that erodes its atmosphere, and determines its ultimate mass and radius over cosmic time scales. Based on our results, HIP 67522  b is continually exposed to a roughly six times higher flare rate than it would be without interaction. Observations with the James Webb Space Telescope suggest that the bulk density of HIP 67522  b is extremely low: Despite its Jupiter-sized radius, its mass was constrained to less than $0.05M_{\rm Jup}$, with a best-fit mass of $15M_{\oplus}$ [7]. The low mass and strong irradiation by its young host star are considered the main reasons for the high inflation of HIP 67522  b’s atmosphere [7, 21]. The total amount of mass lost in the upcoming 100 million years will determine whether HIP 67522  b becomes a hot Neptune type planet, or will further erode and lose a significant fraction of its atmosphere to become a smaller, sub-Neptune object [7]. Our results suggest that previous estimates, although already indicating considerable mass loss [7, 21] from high energy radiation, may still be underestimating the total mass loss rate that can occur in HIP 67522  b under the exposure to the elevated flare rate. On the Sun, large flares are associated with coronal mass ejections (CMEs). On other stars, a CME is estimated to carry roughly the same the energy as the flare itself [22]. If a CME is launched from the footpoint of flaring interaction near the sub-planetary point, it is much more likely to hit the planet than if the CME was triggered in random locations on the star. Under energy-limited escape, with a typical heating efficiency of the planet atmosphere of $10\%$ and CME angular widths typically observed from the Sun ($50^{\circ}-80^{\circ}$ [23]), the impact of planet-induced CMEs would increase the mass loss rate derived from quiescent X-ray und UV flux [7] by $50-130\%$ (Extended Data Fig. 10). Consequently, self-induced flaring may considerably shorten the lifetime of the planet’s atmosphere from $1000$ to about $400-700$ million years.
@@ Methods
We removed the variability from the TESS light curves using the approach detailed in [10]. The CHEOPS data could not be treated in the same way because of the much shorter observing baselines and frequent gaps in the observations. We modeled each $13-23\,$h CHEOPS visit individually with a combination of the batman [25] transit model for HIP 67522  b with quadratic limb darkening using the planetary orbital and limb darkening parameters from [6], and a 5th degree polynomial to approximate the star’s rotational variability. Three large flares and all data points with a CHEOPS quality flag $>0$ were masked prior to fitting the model using a least-square fit (scipy.curve_fit[26]). We then subtracted the model from the data, and masked all outliers above 4 standard deviations in the residuals. We then repeated the model fit with the updated mask to obtain the final model. The model was then subtracted again from the data, and we applied a Savitzy-Golay [27] filter to smooth the remaining non-flaring variability. For the filter, we chose a 3rd order polynomial, and a window that was $20\%$ of the total length of each visit with CHEOPS. As a final correction, we subtracted the remaining roll dependent systematics that were not captured by PIPE using the median of the 100 closest data points in roll angle as the truth for any data point in the light curve.
For the initial flare identification from the resulting flattened light curves, we followed the approach in [10], where three consecutive data points three standard deviations above the noise level define a candidate event. The noise levels of the CHEOPS and TESS light curves are similar, $810\pm 90\,$ppm and $780\pm 70\,$ppm at 101010\,10s and 120120120\,120s cadence, respectively, defined as the standard deviation of the de-trended light curves.
@@
We calculated the bolometric energy of the flares following the prescription in [30], which converts the flux emitted by the star and the flare in the TESS and CHEOPS passbands to the bolometric emission from the flare. We assumed the temperatures of both the star and the flare as known. We propagated the posterior distribution of the flare parameters to derive the bolometric flare energy using Gaussian distributions for the stellar radius $R_{*}=1.39\pm 0.06$ [5] and effective temperature $T_{\rm eff}=5650\pm 75$ [5]. We assumed that the flares were black body emitters with a temperature of $10^{4}\,$. We note that, on the one hand, the flare temperature can vary greatly [31, 32]. On the other hand, the flare energy is not very sensitive to the assumed flare temperature when using the prescription in [30].
@@
assume that $\lambda_{0}$ and $\lambda_{1}$ are independent, and take values between 00 and $2$. We did not make assumptions about where the clustering begins ($\phi_{0}$) or how wide the clustering phase range is ($\Delta\phi$). Therefore, we let $\phi_{0}$ run from 0 to 1 with equal probability. We restrict $\Delta\phi$ from 0 to 0.5 so that we avoid double-counting the identical solution pairs generated by the substitutions: $\phi_{0}\rightarrow\phi_{0}+\Delta\phi$, $\Delta\phi\rightarrow 1-\Delta\phi$, $\lambda_{0}\rightarrow\lambda_{1}$ and $\lambda_{1}\rightarrow\lambda_{0}$ (Note that we do not impose $\lambda_{1}>\lambda_{0}$). Normalization of the prior probability of $\Delta\phi$ contributes the factor of $2$ to the prior. The marginal likelihood for $M_{\text{mod}}$ is:
@@
In the unmodulated flaring model, the best-fit flare rate is $\lambda_{0}=0.21^{+0.06}_{-0.05}\,$d-1above the minimum detected flare energy of $1.0\times 10^{34}\,$erg, consistent with the simple division of 15 flares by the total observing time of 73.273.273.2\,73.2d. For the modulated flaring model, the flare rate outside the clustering range is lower, $\lambda_{0}=0.09^{+0.06}_{-0.04}\,$d-1. The planet-induced flaring rate is roughly six times higher than the base rate, at $\lambda_{1}=0.59^{+0.38}_{-0.22}\,$d-1. The flare rate is elevated starting at around transit, i.e., $\phi_{0}=0.00^{+0.02}_{-0.08}$, and lasts for about $20\%$ of the orbit, i.e., $\Delta\phi=0.20^{+0.11}_{-0.12}$.
We calculated the flux emitted in flares through planet-induced flaring by extrapolating the flare frequency distribution to lower and higher flare energies. We used the difference in flare rate $\lambda_{1}-\lambda_{0}$ and the power law slope $\alpha\approx 1.6$ derived from the flare frequency distribution of the full sample of 15 flares following the method in [34]. Since the power law slope is $\alpha<2$, the total flux is dominated by high energy flares. For a conservative estimate, we set the maximum energy of planet-induced flares to the highest flare energy detected in the elevated flaring range, that is $7.4\times 10^{35}\,$erg. Flares at and above this energy occur on a monthly basis. With the chosen maximum energy, the total flux in planet-induced flares is $4.6\times 10^{29}\,$erg/s. Even if the three highest energy events out of the 11 flares in the elevated flaring region were not planet-induced but stemmed from the intrinsic $\lambda_{0}$ contribution, the planet-induced flux would still be $>10^{29}\,$erg/s
The planet-induced flare flux is considerably higher than the interaction flux expected in the Alfvén wing model [1], even under very generous assumptions. Using the same procedure as in [10], we calculate the average magnetic field strength of HIP 67522 from its X-ray luminosity from [21], and alternatively, from its Rossby number, following [35], which yield $3.1$ and 2.12.12.1\,2.1kG, respectively. For planetary magnetic fields $B_{p}=0.01-100\,$G, stellar wind densities at the coronal base of $\rho=0.1-100\rho_{\odot}$ ($\rho_{\odot}=10^{10}\,$cm-3), and stellar average magnetic field strengths $B=1600-3500\,$G, the star-planet interaction flux yields $10^{24}-10^{28}\,$erg/s, where the strongest flux is obtained with the highest values in the given ranges. Therefore, it is unlikely that the Alfvén wing interaction alone powers the elevated flaring in HIP 67522 .
@@
Table 2: Orbital phase coverage.
@@
22footnotetext: The unbinned phase coverage table is provided online.
Table 3: TESS and CHEOPS flares.
@@
CHEOPS 2460392.789 0.048 35.9+${}_{-0.04}^{+0.03}$ 0.025884
CHEOPS 2460400.003 0.012 34.9+${}_{-0.04}^{+0.04}$ 0.062428
CHEOPS 2460455.637 0.003 34.1+${}_{-0.12}^{+0.09}$ 0.056482
CHEOPS 2460413.181 0.002 33.3+${}_{-0.08}^{+0.07}$ 0.955996 excluded1
TESS 2458608.290 0.005 34.6+${}_{-0.05}^{+0.04}$ 0.612894
TESS 2458609.676 0.011 34.7+${}_{-0.04}^{+0.04}$ 0.812154
TESS 2459334.890 0.010 35.2+${}_{-0.08}^{+0.07}$ 0.017420
TESS 2459334.981 0.024 35.5+${}_{-0.05}^{+0.05}$ 0.030477
TESS 2459335.173 0.006 34.9+${}_{-0.14}^{+0.10}$ 0.058039
TESS 2459342.014 0.008 34.6+${}_{-0.05}^{+0.04}$ 0.041006
TESS 2460042.757 0.005 34.3+${}_{-0.07}^{+0.06}$ 0.730162
TESS 2460044.282 0.005 34.5+${}_{-0.05}^{+0.05}$ 0.949207
TESS 2460045.424 0.003 34.0+${}_{-0.11}^{+0.09}$ 0.113382
TESS 2460052.991 0.004 34.3+${}_{-0.06}^{+0.05}$ 0.200551
TESS 2460059.608 0.049 35.5+${}_{-0.04}^{+0.04}$ 0.151479
TESS 2460065.829 0.004 34.2+${}_{-0.07}^{+0.06}$ 0.045315
@@
11footnotetext: Flare was excluded from the statistical analysis because it was below the detection threshold of TESS.
@@
Figure 11: Flare frequency distributions of G dwarf stars with rotation periods $0.27-5.41\,$d, and effective temperatures $5113-5916\,$K with at least three flares observed with TESS [33], excluding known eclipsing binaries from [36]. We set the upper limit on the flare rates $\lambda_{0}$ and $\lambda_{1}$ conservatively to 222\,2d-1 (horizontal dashed line) above the minimum flare energy in the sample (vertical dashed line). Note that the [33] sample still likely contains binary stars, such that the true flare rate of some stars is lower than shown.

### v2

@@ Abstract
HIP 67522 is a young, 17 million year old star-planet system in the Upper Centaurus Lupus part of the Sco-Cen OB association [8, 5] about 125125\,pc away [9]. The star is a $1.2M_{\odot}$ dwarf [5] that hosts one gas giant planet in a 6.956.95\,d orbit [5], and another in a 14.3314.33\,d orbit [6].
@@
HIP 67522 was previously observed by the Transiting Exoplanet Survey Satellite (TESS [15]) in Sectors 11 (May 2019), 38 (May 2021), and 64 (April 2023) at a 2-min cadence for a total of 65.665.6\,d (Extended Data Fig. 5). Additionally, we gathered dedicated observations of the HIP 67522 system with the CHaracterising ExoPlanets Telescope (CHEOPS [16]) for a total of 7.67.6\,d between March 9 and June 22, 2024 (Extended Data Table 1). We achieved excellent phase coverage: each hour of the 167167\,h orbit of HIP 67522  b was fully covered at least seven times (Fig. 3 and Extended Data Table 2). We detected 12 flares in the TESS light curves (Extended Data Fig. 6), and 4 in the CHEOPS light curves (Extended Data Fig. 7, see also Methods). Of the CHEOPS flares, 3 were above the detection threshold for TESS. We use these for analysis, giving a total of 15 flares (Extended Data Table 3), to ensure that the different detection thresholds for flares in CHEOPS and TESS do not bias our conclusion.
Figure 2: Observed flare rates on HIP 67522 relative to the orbit of the innermost planet HIP 67522  b. HIP 67522 showed a significant increase in flare rate in the expected range, indicating that the majority of detected flares were triggered by the planet. The orange dashed region illustrates the expected rate of planet-induced flares from a footpoint of interaction at the sub-planetary longitude. Planet-induced flares can only be seen when the planet faces the observer ($\pm 90^{\circ}$), and most flares can be expected near the transit of HIP 67522  b ($0^{\circ}$).
@@
The best-fit center of the clustering region is at phase $0.08\pm 0.04$, consistent with being centered on transit within $2\sigma$. The slight offset from transit may therefore be coincidental. Alternatively, it may be due to a distinctive feature of the HIP 67522 system. With a rotation period of approximately 1.421.42\,d, HIP 67522  b orbits near a 1:5 commensurability between the rotation period of the star and the orbital period of HIP 67522  b. If this resonance was exact, the planet could align with the same stellar longitude at fixed orbital phases. In this scenario, a misaligned stellar magnetic field (Fig. 1) would cause the planet to encounter periodic maxima and minima in magnetic field strength multiple times per orbit. This varying field induces harmonic modulation in the star-planet interaction strength, which depends on the position of the footpoint, and the stellar magnetic field strength encountered by the planet [1, 19]. This harmonic modulation introduces two parameters into the model: the dipole field obliquity $\theta$ and the initial dipole azimuthal orientation $\phi_{0}$ relative to the observer. These parameters reproduce the peak power of interaction in the observed phase range for about $17\%$ of all possible configurations, spanning all obliquities above $20^{\circ}$ (Fig. 4). If the system is not in resonance, this modulation would average out over time, leaving the clustering determined only by self-shadowing and visibility of the footpoint, similar to Fig. 2.
@@
Regardless of the underlying mechanism, the observed planet-induced emission has striking consequences. The orbital phases of the flares induced by HIP 67522  b imply that the interaction with the star backfires on the planet through self-inflicted space weather. A planet’s space weather denotes the bombardment with high energy radiation and particles that erodes its atmosphere, and determines its ultimate mass and radius over cosmic time scales. Based on our results, HIP 67522  b is continually exposed to a roughly six times higher flare rate than it would be without interaction. Observations with the James Webb Space Telescope suggest that the bulk density of HIP 67522  b is extremely low: Despite its Jupiter-sized radius, its mass was constrained to less than $0.05M_{\rm Jup}$, with a best-fit mass of $15M_{\oplus}$ [7]. The low mass and strong irradiation by its young host star are considered the main reasons for the high inflation of HIP 67522  b’s atmosphere [7, 21]. The total amount of mass lost in the upcoming 100 million years will determine whether HIP 67522  b becomes a hot Neptune type planet, or will further erode and lose a significant fraction of its atmosphere to become a smaller, sub-Neptune object [7]. Our results suggest that previous estimates, although already indicating considerable mass loss [7, 21] from high energy radiation, may still be underestimating the total mass loss rate that can occur in HIP 67522  b under the exposure to the elevated flare rate. On the Sun, large flares are associated with coronal mass ejections (CMEs). On other stars, a CME is estimated to carry roughly the same the energy as the flare itself [22]. If a CME is launched from the footpoint of flaring interaction near the sub-planetary point, it is much more likely to hit the planet than if the CME was triggered in random locations on the star. Under energy-limited escape, with a typical heating efficiency of the planet atmosphere of $10\%$ and CME angular widths typically observed from the Sun ($50^{\circ}-80^{\circ}$ [23]), the impact of planet-induced CMEs would increase the mass loss rate derived from quiescent X-ray und UV flux [7] by $50-130\%$ (Extended Data Fig. 10). Consequently, self-induced flaring may considerably shorten the lifetime of the planet’s atmosphere from $1000$ to about 400700400-700 million years.
@@ Methods
We removed the variability from the TESS light curves using the approach detailed in [10]. The CHEOPS data could not be treated in the same way because of the much shorter observing baselines and frequent gaps in the observations. We modeled each 132313-23\,h CHEOPS visit individually with a combination of the batman [25] transit model for HIP 67522  b with quadratic limb darkening using the planetary orbital and limb darkening parameters from [6], and a 5th degree polynomial to approximate the star’s rotational variability. Three large flares and all data points with a CHEOPS quality flag >0>0 were masked prior to fitting the model using a least-square fit (scipy.curve_fit[26]). We then subtracted the model from the data, and masked all outliers above 4 standard deviations in the residuals. We then repeated the model fit with the updated mask to obtain the final model. The model was then subtracted again from the data, and we applied a Savitzy-Golay [27] filter to smooth the remaining non-flaring variability. For the filter, we chose a 3rd order polynomial, and a window that was $20\%$ of the total length of each visit with CHEOPS. As a final correction, we subtracted the remaining roll dependent systematics that were not captured by PIPE using the median of the 100 closest data points in roll angle as the truth for any data point in the light curve.
For the initial flare identification from the resulting flattened light curves, we followed the approach in [10], where three consecutive data points three standard deviations above the noise level define a candidate event. The noise levels of the CHEOPS and TESS light curves are similar, $810\pm 90\,$ppm and $780\pm 70\,$ppm at 1010\,s and 120120\,s cadence, respectively, defined as the standard deviation of the de-trended light curves.
@@
We calculated the bolometric energy of the flares following the prescription in [30], which converts the flux emitted by the star and the flare in the TESS and CHEOPS passbands to the bolometric emission from the flare. We assumed the temperatures of both the star and the flare as known. We propagated the posterior distribution of the flare parameters to derive the bolometric flare energy using Gaussian distributions for the stellar radius $R_{*}=1.39\pm 0.06$ [5] and effective temperature $T_{\rm eff}=5650\pm 75$ [5]. We assumed that the flares were black body emitters with a temperature of $10^{4}\,$K. We note that, on the one hand, the flare temperature can vary greatly [31, 32]. On the other hand, the flare energy is not very sensitive to the assumed flare temperature when using the prescription in [30].
@@
assume that $\lambda_{0}$ and $\lambda_{1}$ are independent, and take values between 0 and $2$. We did not make assumptions about where the clustering begins ($\phi_{0}$) or how wide the clustering phase range is ($\Delta\phi$). Therefore, we let $\phi_{0}$ run from 0 to 1 with equal probability. We restrict $\Delta\phi$ from 0 to 0.5 so that we avoid double-counting the identical solution pairs generated by the substitutions: $\phi_{0}\rightarrow\phi_{0}+\Delta\phi$, $\Delta\phi\rightarrow 1-\Delta\phi$, $\lambda_{0}\rightarrow\lambda_{1}$ and $\lambda_{1}\rightarrow\lambda_{0}$ (Note that we do not impose $\lambda_{1}>\lambda_{0}$). Normalization of the prior probability of $\Delta\phi$ contributes the factor of $2$ to the prior. The marginal likelihood for $M_{\text{mod}}$ is:
@@
In the unmodulated flaring model, the best-fit flare rate is $\lambda_{0}=0.21^{+0.06}_{-0.05}\,$d-1above the minimum detected flare energy of $1.0\times 10^{34}\,$erg, consistent with the simple division of 15 flares by the total observing time of 73.273.2\,d. For the modulated flaring model, the flare rate outside the clustering range is lower, $\lambda_{0}=0.09^{+0.06}_{-0.04}\,$d-1. The planet-induced flaring rate is roughly six times higher than the base rate, at $\lambda_{1}=0.59^{+0.38}_{-0.22}\,$d-1. The flare rate is elevated starting at around transit, i.e., $\phi_{0}=0.00^{+0.02}_{-0.08}$, and lasts for about $20\%$ of the orbit, i.e., $\Delta\phi=0.20^{+0.11}_{-0.12}$.
We calculated the flux emitted in flares through planet-induced flaring by extrapolating the flare frequency distribution to lower and higher flare energies. We used the difference in flare rate $\lambda_{1}-\lambda_{0}$ and the power law slope $\alpha\approx 1.6$ derived from the flare frequency distribution of the full sample of 15 flares following the method in [34]. Since the power law slope is $\alpha<2$, the total flux is dominated by high energy flares. For a conservative estimate, we set the maximum energy of planet-induced flares to the highest flare energy detected in the elevated flaring range, that is $7.4\times 10^{35}\,$erg. Flares at and above this energy occur on a monthly basis. With the chosen maximum energy, the total flux in planet-induced flares is $4.6\times 10^{29}\,$erg/s. Even if the three highest energy events out of the 11 flares in the elevated flaring region were not planet-induced but stemmed from the intrinsic $\lambda_{0}$ contribution, the planet-induced flux would still be $10^{29}\,$erg/s
The planet-induced flare flux is considerably higher than the interaction flux expected in the Alfvén wing model [1], even under very generous assumptions. Using the same procedure as in [10], we calculate the average magnetic field strength of HIP 67522 from its X-ray luminosity from [21], and alternatively, from its Rossby number, following [35], which yield $3.1$ and 2.12.1\,kG, respectively. For planetary magnetic fields $B_{p}=0.01-100\,$G, stellar wind densities at the coronal base of $\rho=0.1-100\rho_{\odot}$ ($\rho_{\odot}=10^{10}\,$cm-3), and stellar average magnetic field strengths B=16003500B=1600-3500\,G, the star-planet interaction flux yields $10^{24}-10^{28}\,$erg/s, where the strongest flux is obtained with the highest values in the given ranges. Therefore, it is unlikely that the Alfvén wing interaction alone powers the elevated flaring in HIP 67522 .
@@
Table 2: Orbital phase coverage.22footnotetext: The unbinned phase coverage table is provided online.
@@
Table 3: TESS and CHEOPS flares.11footnotetext: Flare was excluded from the statistical analysis because it was below the detection threshold of TESS.
@@
[…]
@@
Figure 11: Flare frequency distributions of G dwarf stars with rotation periods 0.275.410.27-5.41\,d, and effective temperatures 511359165113-5916\,K with at least three flares observed with TESS [33], excluding known eclipsing binaries from [36]. We set the upper limit on the flare rates $\lambda_{0}$ and $\lambda_{1}$ conservatively to 22\,d-1 (horizontal dashed line) above the minimum flare energy in the sample (vertical dashed line). Note that the [33] sample still likely contains binary stars, such that the true flare rate of some stars is lower than shown.